# The Unreasonable Effectiveness of Easy Training Data for Hard Tasks

### v1

@@ 7 Conclusion
Acknowledgements
We thank Yanai Elazar, Kyle Richardson, Ben Bogin, and Tushar Khot for helpful feedback on this paper, as well as Joris Baan, Nouha Dziri, and others at the Allen Institute for AI for useful conversations about the work. We especially thank Oyvind Tafjord for assistance regarding the human hardness measures we use. This work was conducted while Peter Hase was a research intern at the Allen Institute for AI.
@@ Appendix B Additional Results
Figure 10: Correlations between hardness measures across datasets. We omit MDL for GSM8k because 7b parameter models obtain extremely high loss on GSM8k problems, and MDL is valid as a metric only when using a reasonably good model of the data.
Figure 11: Test accuracy declines as test data hardness increases (shown for additional hardness measures).
Figure 12: Test accuracy declines as test data hardness increases (shown for QLoRA with Llama-2-70b).
Figure 13: Easy-to-hard generalization measured on all test data (not subsetting to hard test data), while training on easy/hard data defined according to each hardness measure. Results are similar to testing on hard data, except for GSM8k, where accuracy on the whole data distribution becomes comparable (training on easy data outperforms hard data on easy/medium test data).
Figure 14: Easy-to-hard generalization measured on easy test data, while training on easy/hard data defined according to each hardness measure. This plot shows hard-to-easy generalization for each dataset, compared to easy-to-easy genearlization. On some datasets, hard data makes for better training data, while for others, easy training data is better for easy test performance.
@@
Figure 16: Easy-to-hard generalization results for different models on StrategyQA. Results are similar for Llama-2-70b and Mixtral, while Qwen appears to have been trained on StrategyQA data in a CoT format.
Figure 17: Test performance on hard data sometimes declines more significantly as the gap between train and test hardness grows, but often the difference between training on Medium and Easy data is relatively small in nature. MMLU not shown here since there are only two hardness levels for that dataset (high school vs. college). See Fig. 9 for more results training on college vs. high school vs. 8th grade vs. 3rd grade data.
@@ Appendix C Dataset Details
We provide additional details for each dataset below, including test data subsetting decisions (see final sample sizes in Table 2). All datasets are publicly available. See also Table 1 for a list of which hardness measures are available for which datasets, as well as sample sizes. See Fig. 20 for histograms of hardness measures per dataset.
@@ Appendix D Modeling and Tuning Details
We provide additional information around model tuning for each training method here. All training methods use the same prompts, one per dataset, that were selected based on their success in past work Saha et al. (2023). We show prompts for ARC and MMLU in Table 3. In this template, the {} placeholders are filled with the question, four answer choices, and a single answer choice (explained next). We use this prompt for multiple choice scoring of the four answer choices for all methods, meaning that we run four forward passes to either (1) compute answer choice probabilities for each answer choice for ICL and QLoRA, or (2) collect final answer choice token representations for each answer choice for linear probing. Thus the final answer slot, “A: {}” is populated by each answer choice once. When prompting with $k$ in-context examples for ICL, we separate examples with a line break (one line between each pair of examples).
@@
For ICL, we select $k=10$ for ARC and MMLU and $k=8$ for StrategyQA and GSM8k as we see diminishing returns to accuracy from larger values of $k$, and using larger $k$ values significantly slows down experiments.
@@
For Linear Probing, we fit a linear classifier to frozen LM hidden states. For a given question, we compute one representation per answer choice by concatenating the question and answer choice and feeding it to the model. To get a single representation from the LM forward pass, we concatenate the representations at the last token index (i.e., the last answer token) from the middlemost and last layer. Then, we score each question-answer choice representation $z$ by applying the linear probe: $f(z;w)=w^{T}z$. The answer choice with the highest score is returned as the prediction. The probe weight $w$ is trained using SGD to minimize cross-entropy loss on a dataset of frozen representations $Z=\{\{z_{i,j}\}_{j=1}^{j=|A|}\}_{i=1}^{N}$ for a dataset of $N$ training datapoints and $|A|$ answer choices. We optimize the weight with the default SGD implementation in PyTorch Paszke et al. (2019) for 100 epochs, without early stopping on any dev data. Based on tuning experiments with Llama-2-13b on ARC, we chose to use SGD rather than AdamW, selecting the number of epochs for convergence, and we chose to use the middlemost and last layer representations (concatenated) rather than either on its own. Note this produces very high dimensional inputs, but by using $\ell_{2}$ regularization with $\lambda=1$, we are able to stably fit probes to these $2d$-dimensional input representations (where $d=8192$ for Llama-2-70b) with as few as $n=10$ training points. The learning rate was fixed at 5e-2.
@@
For QLoRA, we selected hyperparameters based on early experiments with Llama-13b on ARC. Based on this setting, we selected an adaptor rank of $r=16$ rather than $r=8$, with default hyperparameters otherwise, including default selected layers to optimize. The learning rate was fixed at 1e-4, and the model is optimized with a batch size of 50. At least 10 gradient updates are performed, or a minimum of three epochs, whichever yields more gradient updates. This means that for train $n=160$ points and a batch size of 50, we generally perform 12 gradient update steps (3 epochs) in our experiments.
@@
To select the number of decoding steps for each datasets ($t=100$ for StrategyQA and $t=300$ for GSM8k), we wanted to make sure that we were generating reasoning chains long enough for Llama-2-70b to solve hard test questions. Therefore, we intentionally selected this parameter based on hard test performance, in order to use as small a valuable as possible (based on experiment efficiency) that did not compromise test performance on hard data.
@@
All models are run in 8bit quantization, except for Qwen and Mixtral, which are run in 16bit format, and falcon-7b and persimmon-8b, which are run in full precision. We observe no performance loss from quantization in any experiment.
@@
Table 4: Model accuracy table when trained on randomly selected data from the whole data distribution and tested on the whole data distribution (zero-shot ICL rows included), using Llama-2-70b. Averaged over 5 seeds. Compare to Fig. 13.

### v2

@@ 7 Conclusion
Limitations
We aim to study how models generalize from settings where humans can easily label data to those where humans have difficulty labeling data. Since we require ground truth labels to evaluate model generalization, this limits our evaluations to datasets where humans have reliably labeled the data. While we aim to test a broad range of train and test difficulties (like 3rd grade training data and college-level test data), our results may not generalize to settings with difficulty levels besides those we test in this paper, especially test settings where almost no amount of human effort can reliably produce accurate data labels (like unsolved scientific questions).
Additionally, we are not able to verify that test questions for ARC, MMLU, StrategyQA, and GSM8k are not in training datasets of the open models used in this paper, including Llama-2-70b and Mixtral 8x-7b. This means we cannot be certain that models are generalizing from easy to hard data, rather than reporting memorized hard question answers. While this concern currently affects all research on open source LLMs, including those with public datasets (since finding test set contamination in pretraining data is an unsolved problem), we note that it would be beneficial for future work to evaluate easy-to-hard generalization on test sets that are either private or collected after pretraining data collection cut-offs.
Ethics Statement
We hope that positive results in easy-to-hard generalization imply that we can train models to perform well in niche domain like biology, chemistry, medicine, law, engineering, etc., without demanding significant amounts of expert time and spending large amounts of money in order to annotate data in these settings. In this way, we might make LLMs more useful for hard tasks while requiring less human effort to supervise their training. At the same time, we note that there are dual use and human labor displacement concerns around improving model capabilities in such domains. Ultimately we hope for LLMs to be deployed responsibly, so that they can be used to further human values and not for any ill intent.
@@ Appendix B Additional Results
    We show easy-to-hard generalization on additional hardness measures in Fig. 18.
  6. 6.
@@
  11. 11.
    In Fig. 19, we test the hypothesis that task format alone is taught by training data, as opposed to true generalization. We list examples used for “task format only” prompts, which are trivially simple examples matching multiple-choice or yes/no answer prompts in our data.
Figure 8: Correlations between hardness measures across datasets (Spearman rank order correlation). We omit MDL for GSM8k because 7b parameter models obtain extremely high loss on GSM8k problems, and MDL is valid as a metric only when using a reasonably good model of the data.
Figure 9: Test accuracy declines as test data hardness increases (shown for additional hardness measures), with the exception of MMLU Answer Chars. Error bars are 95% CIs showing test sample variance.
Figure 10: Test accuracy declines as test data hardness increases (shown for QLoRA with Llama-2-70b). Error bars are 95% CIs showing test sample variance.
Figure 11: Easy-to-hard generalization measured on all test data (not subsetting to hard test data), while training on easy/hard data defined according to each hardness measure (using Llama-2-70b prompted with $k\leq 10$ examples). Results are similar to testing on hard data, except for GSM8k, where accuracy on the whole data distribution becomes comparable (training on easy data outperforms hard data on easy/medium test data).
Figure 12: Easy-to-hard generalization measured on easy test data, while training on easy/hard data defined according to each hardness measure (using Llama-2-70b prompted with $k\leq 10$ examples). This plot shows hard-to-easy generalization for each dataset, compared to easy-to-easy genearlization. On some datasets, hard data makes for better training data, while for others, easy training data is better for easy test performance.
@@
Figure 14: Easy-to-hard generalization results for different models on StrategyQA, using Llama-2-70b prompted with $k=4$ examples using CoT. Results are similar for Llama-2-70b and Mixtral, while Qwen appears to have been trained on StrategyQA data in a CoT format.
Figure 16: Test performance on hard data sometimes declines more significantly as the gap between train and test hardness grows, but often the difference between training on Medium and Easy data is relatively small in nature (using Llama-2-70b prompted with $k\leq 10$ examples). MMLU not shown here since there are only two hardness levels for that dataset (high school vs. college). See Fig. 7 for more results training on college vs. high school vs. 8th grade vs. 3rd grade data.
@@ Appendix C Dataset Details
We provide additional details for each dataset below, including test data subsetting decisions (see final sample sizes in Table 2). All datasets are publicly available, and license information is included via the links provided. All datapoints are in English. See also Table 1 for a list of which hardness measures are available for which datasets, as well as sample sizes. See Fig. 20 for histograms of hardness measures per dataset.
@@ Appendix D Modeling and Tuning Details
We provide additional information around model tuning for each training method here.
GPU Cost. We run experiments on NVIDIA A6000 GPUs. A typical experiment setting is running Llama-2-70b over $n=2000$ datapoints, quantized 8bit, with ICL with $k=8$, with batch size 1, using 5 random seeds for prompt data selection. Such an experiment requires 4 GPUs and takes about 10 hours to complete. Experiment with CoT are more expensive, taking about one hour per $n=200$ test points on GSM8k, using a decoding length of $t=300$ tokens. Linear probing experiments have negligble runtime as we save the hidden states to file, avoiding the need to rerun model forward passes, while QLoRA experiments take about as long as ICL experiments (slightly faster for CoT settings).
Prompt Templates. All training methods use the same prompts, one per dataset, that were selected based on their success in past work Saha et al. (2023). We show prompts for ARC and MMLU in Table 3. In this template, the {} placeholders are filled with the question, four answer choices, and a single answer choice (explained next). We use this prompt for multiple choice scoring of the four answer choices for all methods, meaning that we run four forward passes to either (1) compute answer choice probabilities for each answer choice for ICL and QLoRA, or (2) collect final answer choice token representations for each answer choice for linear probing. Thus the final answer slot, “A: {}” is populated by each answer choice once. When prompting with $k$ in-context examples for ICL, we separate examples with a line break (one line between each pair of examples).
@@
In-context Learning. For ICL, we select $k=10$ for ARC and MMLU and $k=8$ for StrategyQA and GSM8k as we see diminishing returns to accuracy from larger values of $k$, and using larger $k$ values significantly slows down experiments.
@@
Linear Probing. For Linear Probing, we fit a linear classifier to frozen LM hidden states. For a given question, we compute one representation per answer choice by concatenating the question and answer choice and feeding it to the model. To get a single representation from the LM forward pass, we concatenate the representations at the last token index (i.e., the last answer token) from the middlemost and last layer. Then, we score each question-answer choice representation $z$ by applying the linear probe: $f(z;w)=w^{T}z$. The answer choice with the highest score is returned as the prediction. The probe weight $w$ is trained using SGD to minimize cross-entropy loss on a dataset of frozen representations $Z=\{\{z_{i,j}\}_{j=1}^{j=|A|}\}_{i=1}^{N}$ for a dataset of $N$ training datapoints and $|A|$ answer choices. We optimize the weight with the default SGD implementation in PyTorch Paszke et al. (2019) for 100 epochs, without early stopping on any dev data. Based on tuning experiments with Llama-2-13b on ARC, we chose to use SGD rather than AdamW, selecting the number of epochs for convergence, and we chose to use the middlemost and last layer representations (concatenated) rather than either on its own. Note this produces very high dimensional inputs, but by using $\ell_{2}$ regularization with $\lambda=1$, we are able to stably fit probes to these $2d$-dimensional input representations (where $d=8192$ for Llama-2-70b) with as few as $n=10$ training points. The learning rate was fixed at 5e-2.
@@
Model Finetuning with QLoRA. For QLoRA, we selected hyperparameters based on early experiments with Llama-13b on ARC. Based on this setting, we selected an adaptor rank of $r=16$ rather than $r=8$, with default hyperparameters otherwise, including default selected layers to optimize. We use AdamW with a a learning rate fixed at 1e-4, and the model is optimized with a batch size of 50. At least 10 gradient updates are performed, or a minimum of three epochs, whichever yields more gradient updates. This means that for train $n=160$ points and a batch size of 50, we generally perform 12 gradient update steps (3 epochs) in our experiments.
@@
Decoding Steps. To select the number of decoding steps for each datasets ($t=100$ for StrategyQA and $t=300$ for GSM8k), we wanted to make sure that we were generating reasoning chains long enough for Llama-2-70b to solve hard test questions. Therefore, we intentionally selected this parameter based on hard test performance, in order to use as small a valuable as possible (based on experiment efficiency) that did not compromise test performance on hard data.
@@
Quantization. All models are run in 8bit quantization, except for Qwen and Mixtral, which are run in 16bit format, and falcon-7b and persimmon-8b, which are run in full precision. We observe no performance loss from quantization in any experiment.
@@
Table 4: Model accuracy table when finetuned on randomly selected data from the whole data distribution and tested on the whole data distribution (zero-shot ICL rows included), using Llama-2-70b. Averaged over 5 seeds. Compare to Fig. 11.
@@ Appendix E Statistical Testing
The one exception to this kind of bootstrap sampling is for results in Fig. 3 and 9, where we only show CIs derived from test sample variance. In these plots, we average over five random training seeds.
MMLU prompt
Question: Which one is usually yellow?
A) Cheese
B) Apple
C) Carrot
D) Bread
Answer: Cheese
Question: What do we use to wash dishes?
A) Broom
B) Shovel
C) Soap
D) Hammer
Answer: Soap
Question: What color is the sky when it’s sunny?
A) Grey
B) Blue
C) Black
D) Orange
Answer: Blue
StrategyQA prompt
Q: Does cheese come from a plant?
A: no
Q: Can you use a broom to sweep the floor?
A: yes
Q: Is the sky green when it’s sunny?
A: no
Table 5: Example questions used for testing model performance based on task-format-only prompts. We use ChatGPT to generate 50 such questions for MMLU and 40 for StrategyQA. These are used for experiments using $k=10$ examples for MMLU and $k=8$ for StrategyQA, averaging over five different prompts (results shown in Fig. 19).